OpenStudy (anonymous):

16^(x^2-7/4x)=4 solve for x

OpenStudy (anonymous):

Just by looking at this one, you can see that the exponent on 16 must be equal to 1/2 because 16^(1/2) = 4. You get the equation x^x-7/4x = 1/2. Factoring you get (x+1/4)*(x-2) = 0 or x = -1/4 or 2